DOUBLE RESONANCE STUDIES OF COLLISIONAL ENERGY TRANSFER IN CH_{2}

\underline{YANGSOO KIM}, ANATOLY V. KOMISSAROV, GREGORY E. HALL, and TREVOR J. SEARS, Chemistry Department, Brookhaven National Laboratory, Upton, NY 11973-5000.

We have studied rotational energy transfer by optical-optical double resonance methods in order to understand the close connection between rotational energy transfer and the mixed-state mechanism of intersystem crossing. Singlet CH_{2} is produced by 308 nm photolysis of CH_{2}CO. After thermalization of the rotational and translational distribution, a bleaching tunable dye laser pulse depletes the population of a selected rotational level of singlet CH_{2} while the population of the same (in saturation recovery mode) or a different (in saturation transfer mode) singlet CH_{2} rotational level is monitored with FM absorption spectroscopy on a different vibrational band of the b̃ − ã system. The bleaching laser is blocked and unblocked to record a reference and a bleached transient waveform. The difference of the two signals as a fraction of the reference signal is the saturation recovery signal, which is found to follow a single exponential decay to a final value closed to half of the initial thermal Boltzmann fraction of the population in the bleached level. The rate coefficients for saturation recovery by collision with CH_{2}CO and He are 1.6 × 10^{−8} and 2.1 × 10^{−9} cm^{3} molec^{−1} s^{−1}, respectively. The saturation transfer kinetics is multi-exponential and depends on which pair of rotational states is saturated and probed.